QUESTION 19 You hold a portfolio of mortgages $100. 115% of these morgages default, you will lose exactly $ will lose exactly S A bank has used $100 million morgages to create the following tranches: If 12% of these morgages default, you 45%) AAA 25%) AA 18% 7% BBB 5% Residual Equit Imagine that you invest your $100 in the BBB tranch·115% of the underlying morgages default, you will lose exactly Hint: Think of the water cascade! . If 12% of the underlying morgages default, you will lose exactly $ Sit back and compare these two investment choices (directly into morgages, versus indirectly via tranches). See how the risk profile is very different? This is what surprised many investors just before the GFC. The stuned faces in the Big Short as well as Margin Call reflect the sudden realization of this underlying risk.
Answer #1

Answer

1.

a. if 5% of mortgage default then loss will be as follows

=(100*45%+100*25%+100*18%+100*7%+100*5%)*5%= $5

b. if 12% of mortgage default then loss will be as follows

=(100*45%+100*25%+100*18%+100*7%+100*5%)*12%= $12

2. if investor invests $100 only in BBB tranches then loss will be

if default is 5% = 100*7%*5%= $.35

if default is 12%= 100*7%*12%=$.84

comparison between mortgage and tranches investments

we can see if we are using mortgage then we are incurring more fluctuations in data because multi- risk factors used in portfolio.

if we are using only BBB tranche then only one risk factor is available .

as we know higher the risk higher the return .